1. derive the reduction formulas for
  2. ∫cosnxdx satisfies nIn=cosn-1xsinx+(n-1)In-2
  3. ∫xnexdx
  4. ∫tannxdx
1
Expert's answer
2021-07-22T12:58:16-0400

Solution

2) "\u222bcos^nxdx"

We have

"I_n=\u222bcos^nxdx"

"cos^n x=cos^{n-1}xcosx"

"\implies I_n=∫cos^{n-1}xcosxdx"

Let's do integration by parts

Let "u=cos^{n-1}x \implies du=(n-1)cos^{n-2}x(-sinx)dx=-(n-1)cos^{n-2}x(sinx)dx"

Let "dv=cosxdx \implies v=sinx"

"I_n=\int udv=uv-\int vdu"

"=cos^{n-1}xsinx-\int sinx(-(n-1)cos^{n-2}x(sinx))dx"

"=cos^{n-1}xsinx+\int sinx(n-1)cos^{n-2}xsinxdx"

"=cos^{n-1}xsinx+\int (n-1)cos^{n-2}xsinxsinxdx"

but "sinxsinx=sin^2x"

"\implies I_n=cos^{n-1}xsinx+\int (n-1)cos^{n-2}xsin^2xdx"

We know that "sin^2x=1-cos^2x"

"\implies I_n=cos^{n-1}xsinx+\int (n-1)cos^{n-2}x(1-cos^2x)dx"

"=cos^{n-1}xsinx+(n-1)\int cos^{n-2}x(1-cos^2x)dx"

"=cos^{n-1}xsinx+(n-1)\int cos^{n-2}xdx-(n-1)\int cos^nxdx"

"=cos^{n-1}xsinx+(n-1)I_{n-2}-(n-1)I_n"

Now, collect the like terms

"\implies I_n+(n-1)I_n=cos^{n-1}xsinx+(n-1)I_{n-2}"

"\implies I_n+nI_n-I_n=cos^{n-1}xsinx+(n-1)I_{n-2}"

"\therefore nI_n=cos^{n-1}xsinx+(n-1)I_{n-2}"


"\implies I_n={cos^{n-1}xsinx+(n-1)I_{n-2}\over n}" for "n\ge 2"

For initial values "(I_2,I_3....)"


3) "\u222bx^ne^xdx"

Let's do integration by parts

Let "u=x^n \implies du=nx^{(n-1)}dx"

"dv=e^xdx \implies v=e^x"

"I_n=\int udv=uv-\int vdu"

"=x^ne^x-\int e^xnx^{(n-1)}dx"

"=x^ne^x-n\int x^{(n-1)}e^xdx"

"=x^ne^x-nI_{n-1}"

"\therefore I_n=x^ne^x-nI_{n-1}" for "n\ge1"

For initial values "(I_1,I_2...)"

4) "\u222btan^nxdx"

We have

"I_n=\u222btan^nxdx"

"=\u222btan^{n-2}xtan^2xdx" but "tan^2x=sec^2x-1"

"=\u222btan^{n-2}x(sec^2x-1)dx"

"=∫tan^{n-2}xsec^2xdx-\int tan^{n-2}xdx"

Let's evaluate "\u222btan^{n-2}xsec^2xdx"

Let "u=tanx \implies du=sec^2xdx"

Rewrite "\u222btan^{n-2}xsec^2xdx" in terms of "u" as follows

"∫tan^{n-2}xsec^2xdx=\int u^{n-2}du={u^{n-1}\over n-1}={tan^{n-1}x\over n-1}"

"\implies I_n={tan^{n-1}x\over n-1}-\int tan^{n-2}xdx"

"\therefore I_n={tan^{n-1}x\over n-1}-I_{n-2}" for "n\ge2"

For initial values "(I_2,I_3...)"